OpenStudy (anonymous):

Write 5/(3+(sqrt)2) in the form a+b(sqrt)2 where a and b are rational numbers. PLS HALP <3

OpenStudy (anonymous):

multiply the equation by (3-sqrt(2))/(3-sqrt(2)) which is unity so it doesn't actually change the equation.

OpenStudy (anonymous):

(5(3-sqrt())/((3+sqrt(2))(3-sqrt(2))=

OpenStudy (anonymous):

(15-5sqrt(2))/(9+2)

OpenStudy (anonymous):

=(15/11) + (5/11)sqrt(2)

OpenStudy (anonymous):

\[5/(3-\sqrt{2)} \times (3+\sqrt{2})/(3+\sqrt{2})\]

OpenStudy (hazza):